DIRC Operations Report
Wednesday 2007/12/3 to Wednesday 2007/8/15
Oncall: Jose Benitez, Rolf Andreassen, Joe Schwiening
Fast Monitoring: Jose Benitez, Rolf Andreassen, Joe Schwiening
PEP peak lumi: ~8.3 x 1033
Run 6 total delivery: ~7.5 fb-1
Run 6 total BaBar recorded: ~.5 fb-1 on the
Y(4S) , ~6.9 fb-1 on the
Y(3S) , 0 fb-1 off-peak data
PEP activities/issues since last DIRC meeting
First collisions seen on Dec 14.
Switched to running on the Y(3S) on Dec 20.
Smooth running over the holidays, although lumi remained at ~6x10^33
Bad weather last week affected the Linac Friday and Saturday.
Currently having problems with a transverse kicker
magnet. Cant raise the LER current past ~1900mA.
Two day ROD starting today to investigate this problem.
BaBar activities/issues since last DIRC meeting
Power dip on Jan 4th morning (due to weather) affected DCH
and EMC.BaBar Recovered quickly.
IR2 was well prepared for the storm last week. No major
New trigger lines installed to increase efficiency for the new
physics at the Y(3S).
DIRC activities/issues since last DIRC meeting
Nitrogen flow values became unstable on Dec 7th. Problem is
due to flaky flow meter electronics on top of the detector. One
controlled access made on Dec 11 by Jerry, Matt and Steve Dardin,
flakiness disappeared after touchin the electronics
board. Another controlled access was made on Dec 14 by Jerry and
Popat, instabilities went away in the same way.
New relay was installed on the chiller power line on Dec. 13, but we still see
glitches. Next thing to replace is the connector on the back of
the chiller. Also a new flow sensor will be
installed which will be used to monitor the chiller status,
because the current one does not work if the chiller is off.
Backgrounds are generally good, except for a few background
splashes last weekend which resulted in current some current
excess alarms and some unusual FEX errors in bin6.
See DIRC logbook for details
Lumi and Background History
- Controlled access scheduled today to investigate the Nitrogen gas
electronics instabilities on top of the detector.
- The connector for the chiller monitoring lines will be
- HVAC currently installing new flow sensor on chiller. We are
on the BCS.
- One noisy channel in sector 0 will be masked next couple of
days if it persists.
| Equipment || # used || Spares |
| DFB || 168
12/17 good + 4 possibly good, 1 bad
| DCC || 12
| HV Boards || 48
9/11 good , 2 broken still in Angels shop
- Rolf has left on January 4.
- Joe took over the on-call,I am carrying the backup.
- Jennifer currently doing training.